Operations Report – February 08th

Crew 221 Operations Report 07Feb2022

SOL: 6

Name of person filing report: Luis D. Monge

Non-nominal systems:

Notes on non-nominal systems:

Generator: off (at 8pm) waiting for SOC to reach 80% to be turned on

Hours run: 11:00

From what time last night: 8:30pm

To what time this morning: 7:30am

List any additional daytime hours when the generator was run : 0

Solar— SOC % (Before generator is run at night) 89%

Diesel Reading – 97%

Station Propane Reading – 60%

Ethanol Free Gasoline – gallons

Water (loft tank): 35 gallons

Water Meter: 148 units

Water (static tank) – 285 gals

Static to Loft Pump used – yes

Water in GreenHab – 227.3 gals

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: no


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used: yes

Hours: (before EVA) : 129.9

Beginning charge: (Before EVA) 100%

Ending charge: (On return from EVA, before recharging) 75%

Currently charging: yes

Opportunity rover used: no

Hours: n/a

Beginning charge: n/a

Ending charge: n/a

Currently charging: yes

Curiosity rover used: yes

Hours: 136.1

Beginning charge: 100%

Ending charge: 72%

Currently charging: yes

Notes on rovers:

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) not used

Reason for use: not used

Oil Added? not used

ATV Fuel Used: Gals not used

# Hours the ATVs were Used today: not used

Notes on ATVs:

HabCar used and why, where? not used

CrewCar used and why, where? not used

General notes and comments: not used

Summary of internet: operational

Summary of suits and radios: operational, lose cable on suit 7 fixed again but it might neeed a more permanent fix for the lose cable. Suit 10 is not charging, voltage at battery reads 5V, battery change is recomended and permission to perfome the replacement is requested.

Summary of Hab operations: nominal.

Summary of GreenHab operations: operational, black corn experiment continues, now watered with 2.5l/day. watering of the plants carried out as per procedure

Summary of ScienceDome operations: 2 incubators runing, microbiology experiments going on and cultures are bieng grown in agar plates. Detection of K, P and N carried out for tree soil samples currenlty being used at the greenhab for the black corn experiment.

Summary of RAM operations: robot assembly is ongoing adn expected to continue, antenna asembly for radio comunication experiment took place today and is expected to continue.

Summary of any observatory issues: nominal

Summary of health and safety issues: strong gas smell reported and gas detector brough into de the hab, it was determined that the gas came from the oven which was not working properly.

Questions, concerns and requests to Mission Support: the oven was reconfigured to work with propane as the previous confifguration was for natural gas, and that was deteremined to be the cuase of the smell reported yesterday.

Copyright © The Mars Society. All rights reserved. | Main Site